OpenStudy (emogirlatemoooocow):

What is the value of x? 3 • (–4) • x = 0 x =

OpenStudy (anonymous):

0

OpenStudy (killer_shadow):

lol

OpenStudy (xmissalycatx):

x would have to be 12 because 3 (-4) is -12! You want it to be the opposite if -12 because it will cancel out and = 0

OpenStudy (emogirlatemoooocow):

so its 12?

OpenStudy (xmissalycatx):

yes

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. (3)(−4)x=0 Step 1: Simplify both sides of the equation. −12x=0 Step 2: Divide both sides by -12. −12x −12 = 0 −12 x=0